Project

General

Profile

NDOS Commissioning Meeting 2011/08/04

Attendance: Jon, Mark Mathis, Minerba, Gary, Chad, Gavin, Rafael, Susan, Mark, Athans, Andrew, Leon

Agenda:

  1. Beam & Detector Update
  2. DataCheck Update
  3. DAQ Update
  4. DCS Update
  5. AOB

Beam & Detector Update

  • Smooth all week, no issues

DataCheck

  • Plan for getting nearline up and running
    • need dedicated machine for this
    • Andrew: DAQ farm node 16 is available
    • Dominic will be available to work on getting the system up and running
  • link to presentation

DCS Update

  • Athans:
    • Work on GUI continues.
    • Scaling of raw values to physics units.
    • Developing alarm system and integrating HV readout into GUI.

DAQ Update

  • Andrew: lots of tests done on the timing system this week; engineers have not been able to reproduce the problem we are seeing. Also, the TDUs have been in sync all week. So it's hard to catch the problem.
  • Leon:
    • waiting for modifications to test-stand to allow long tests
    • Gary: were any of the wet APDs recovered? Leon: 20 of wet APDs passed initial tests. These also needed to be cleaned. But we won't really know until we do the extended tests. Some "dirty" APDs were recovered by simply blowing dry N2 onto them.
    • power supply for TECC cooler on test-stand
    • Jon: did the env. chamber get fixed? Leon: yes, it's now operational. Tests have been done to try to measure humidity levels inside seal, but none have shown anything above 5%.

AOB

  • Jon has made some changes to the wiki and ECL forms that hopefully are an improvement. Feedback is always welcome.